OpenStudy (anonymous):

(sqrt)5 + 3(sqrt)10 + 2 (sqrt)5

OpenStudy (lgbasallote):

3 sqrt 5 + 3 sqrt 10 so 3( sqrt 5 + sqrt 10)

OpenStudy (anonymous):

\[\sqrt{5} + 3\sqrt{10} +2\sqrt{5}\]

OpenStudy (anonymous):

\[ \sqrt{5} + 3\sqrt{10} +2\sqrt{5} = 3 \left(\sqrt{5}+\sqrt{10}\right) \]